OpenStudy (anonymous):

Does anyone know the equation for combined variation? Here are the answer choices:

OpenStudy (anonymous):

A. \[z=kxy\] B. \[y=\frac{ k }{ x }\] C. \[z=\frac{ ky }{ x }\] D. \[y=kx\]

hartnn (hartnn):

you still need help ?

hartnn (hartnn):

combined variation is when you have both direct and inverse relation like z= ky and z=k/x both giving you z =ky/x